# Let Me Teach You: Pedagogical Foundations of Feedback for Language Models

### v2

@@ A.1 Task
Understanding tha answer is fairly straightforward – a task has a closed-answer if there is a finite set of correct answers, and an open-answer otherwise. Notably, tasks can contain both elements. For example the task "Write a quality 4-paragraph short-story" has both open- and closed-answer elements. There is no finite set of answer of what a quality story is, but whether a story has 4 paragraphs, or not, is a binary closed-answer task.
@@ B.8 Style
Below we present examples of feedback with different style for Model Answer B. Many variations are possible, and we expect the model’s answer to reflect the style used in some linguistic artifacts (e.g.,, if a particularly informal tone was used, we expect the models’ answer to lean toward lower formality as well).

### v3

@@ 6 Conclusion
Acknowledgements
We gratefully acknowledge the support of the Swiss National Science Foundation (No. 215390), Innosuisse (PFFS-21-29), the EPFL Science Seed Fund, the EPFL Center for Imaging, Sony Group Corporation, and the Allen Institute for AI.
@@ A.1 Task
Understanding the answer is fairly straightforward – a task has a closed-answer if there is a finite set of correct answers, and an open-answer otherwise. Notably, tasks can contain both elements. For example the task "Write a quality 4-paragraph short-story" has both open- and closed-answer elements. There is no finite set of answer of what a quality story is, but whether a story has 4 paragraphs, or not, is a binary closed-answer task.
@@ B.8 Style
Below we present examples of feedback with different style for Model Answer B. Many variations are possible, and we expect the model’s answer to reflect the style used in some linguistic artifacts (e.g., if a particularly informal tone was used, we expect the models’ answer to lean toward lower formality as well).